# Diversity as a Reward: Fine-Tuning LLMs on a Mixture of Domain-Undetermined Data

### v2

@@ B.3 Benchmarks
B.4 Training and Evaluation Details
@@
We implement our approaches using PyTorch [31] v2.4.1, coupled with PEFT v0.12.0 and the Transformers library [42] v4.45.2. Experiments are conducted on a computing platform equipped with four NVIDIA A100 GPUs (40GB), with pre-trained LLMs loaded as 16-bit floating-point numbers. The specific data-model development processes are completed in Data-Juicer Sandbox [9], via integration with the ms-swift [54] training repository, and the OpenCompass [15] evaluation repository.
@@
Appendix C Implementation of DaaR
C.1 Implementation of the Synthetic Data Generation
For implementation details of DaaR’s hyper-parameters, we set the number of seed samples $\mathcal{S}_{k}^{(0)}$ as 5 (ablation in Appendix G.3), the number of anchors in sliding window as 3 (ablation in Appendix G.5), the similarity threshold $\tau$ as 0.9 for Mathematics and 0.85 for others (ablation in Appendix G.6), the terminated iteration as 30 (ablation in Appendix G.4), the hidden layer $\mathcal{M}(\tilde{x})$ as layer-3 (ablation in Appendix G.2).
C.2 Implementation of Model-Aware Training Data Generation
Seed Generation
This phase aims to provide zero-shot LLM-generated seed data that reflects how LLMs inherently perceive the characteristics of domain-specific data. We employ the LLM itself to generate concise domain descriptions across diverse fields illustrated in Prompt C.2, and the prompt shown in Prompt C.2. Due to the difficulty of pre-trained LLMs in providing accurate and coherent responses, we utilize their corresponding Instruct versions for centroid data generation. Specifically, we apply Llama3.1-8B-Instruct for Llama3.1-8B, Qwen2-7B-Instruct for Qwen2-7B, and Qwen2.5-7B-Instruct for Qwen2.5-7B.
Diversity Augmentation
While seed data effectively captures the intrinsic domain knowledge of LLMs, its limited quantity and the constrained diversity of single-prompt generation methods necessitate enhancement. We address these limitations through two complementary strategies including a sliding window mechanism and a rule-based filter. The implementation details are formalized in Prompt C.2 and Eq. (7).
Domain-Aware Clustering
[…]
Reasoning: Reasoning involves the ability to think logically about a situation or problem, to draw conclusions from available information, and to apply knowledge in new situations.
Mathematics: Mathematical skills include the ability to perform calculations, understand mathematical concepts, solve hard and professional math problems, and apply mathematical reasoning.
Coding: Design and generate specific code programs, or apply algorithms and data structures, with code generation in the Output.
[…]
Generate 5 different instruction pairs related to this field with various lengths. Maintain the index format: Instruction [1 … 5].
The response should include three parts:
2. Input: Any information provided to help it understand the instruction. If there is no need to generate, just keep it empty.
3. Output: The expected answer or action.
Keep the generated content focused on {selected_domain}. And do not involve {unselected_domains} related knowledge.
[…]
Output: The expected answer or action.
Keep the generated content focused on {selected_domain}. Do not involve {unselected_domain} related knowledge.
Note that you should generate content strongly unrelated and different to these examples to ensure diversity in the generated output:
Counterexample: {}
Appendix D Details of Baselines
D.1 Implementation of Baselines
@@
Random Selection(Rand)
A random selection of 8,000 data samples from the data pool was made, which to some extent reflects the distribution characteristics of the original data pool.
Instruction Length (IL)
The length of the instruction can be considered a measure of input complexity. It is widely believed [5, 53] that more complex data is beneficial for enhancing model capabilities. Therefore, we select a subset of 8,000 entries with the maximum number of words (based on spaces) in the concatenation of Instruction and Input as part of the filtering process.
AlpaGasus [10]
Using prompts to directly score and annotate the quality of the data leverages the powerful cognitive abilities of LLMs for evaluation and selection. Based on the original work, we use the GPT-3.5-Turbo API to score the data with the following prompt:
[…]
@@
Instag [30]
first utilizes ChatGPT to tag the samples based on semantics and intentions, then trains a LLaMA-based tagger on the ChatGPT tags to tag data. They use the number of tags as a proxy for complexity. We directly use ChatGPT-3.5-Turbo as a tagger to achieve better performance. Following the original paper, the prompt is as follows.
[…]
@@
SuperFilter [26]
introduces a method that astonishingly utilizes a small GPT2 model to successfully filter out the high-quality subset from the existing GPT4-generated instruction tuning dataset. The core concept is the instruction-following difficulty (IFD) score. By meticulously following this methodology, we utilize GPT-2 to select 8,000 data samples.
Deita [29]
is an open-sourced project designed to facilitate Automatic Data Selection for instruction tuning in LLMs. It delves into the relationships between data complexity, quality, and diversity, and develops a series of methodologies. Specifically, we utilize it as the following three baselines:
@@
D.2 Visualization of Baseline’s Selected Data
@@
Appendix E Derivation of Proposition 3.2
Expand $q(\mathbf{x})$ via the law of total probability over domains.
$=\frac{\sum_{k=1}^{K}q(c=k)p(\mathbf{x}|c=k)}{p(\mathbf{x})}$ (11)
$=\frac{\sum_{k=1}^{K}q(c=k)\frac{p(c=k|\mathbf{x})p(\mathbf{x})}{p(c=k)}}{p(\mathbf{x})}$ (12)
$=\sum_{k=1}^{K}\frac{q(c=k)}{p(c=k)}p(c=k|\mathbf{x})\cdot\frac{p(\mathbf{x})}{p(\mathbf{x})}$ (13)
Factor out $p(\mathbf{x})$ from the numerator and the denominator.
$=\sum_{k=1}^{K}\underbrace{\frac{q(c=k)}{p(c=k)}}_{\lambda_{k}}p(c=k|\mathbf{x})$ (14)
Define domain proportion ratio $\lambda_{k}:=\frac{q(c=k)}{p(c=k)}$.
@@ F.2 Validation on Qwen3-8B
Following the identical experimental setup outlined in Appendix B, we conducted experiments on Qwen3, with the results presented in Table 9. As demonstrated, DaaR achieved the top comprehensive score on Qwen3, surpassing all baselines by up to 4.7%. Notably, its performance on the MATH benchmark was particularly outstanding, exceeding the average of baselines by 25.91%. The comparative experiments affirm that DaaR maintains competitive effectiveness even on the SOTA LLM Qwen3-8B.
@@ F.3 Customized Data Selection Capability of DaaR
F.4 Rationale for Entropy over Diversity as the Reward Signal
@@ G.6 Ablation of the Diversity Threshold
G.7 Cost-Efficiency and Flexibility
@@ Appendix H Comprehensive Visualization and Experimental Results
Figure 16: The t-SNE visualization of embeddings for data samples with different distributions on Qwen2-7B. (a) The data pool of all 40K samples, (b) Randomly selected subset, (c) Distribution of data farthest from other domain centroids on Inter-Diversity, (d) Distribution of data closest to other domain centroids on Inter-Diversity, (e) Distribution of data closest to its own domain centroid on Inter-Diversity, (f) Distribution of data farthest from its own domain centroid on Inter-Diversity.
Figure 17: The t-SNE visualization of embeddings for data samples with different distributions on Qwen2.5-7B. (a) The data pool of all 40K samples, (b) Randomly selected subset, (c) Distribution of data farthest from other domain centroids on Inter-Diversity, (d) Distribution of data closest to other domain centroids on Inter-Diversity, (e) Distribution of data closest to its own domain centroid on Inter-Diversity, (f) Distribution of data farthest from its own domain centroid on Inter-Diversity.
@@ H.2 Visualization of Inter- & Intra-Diversity Distribution Data
Using the Qwen2-7B model as an example, we construct data based on the Inter-Diversity and Intra-Diversity distribution methods, selecting a batch of data every 20%. The visualization process is shown in Fig. 18 and Fig. 19. As seen in the figures, the data gradually transitions from domain-aware diverse to domain-aware closed, indicating that our data construction method effectively controls the distribution of different data. However, it is important to note that t-SNE dimensionality reduction is used only for visualization purposes; during the reduction process, positional information is lost, and it cannot fully capture the true high-dimensional relationships between the data.
@@ H.3 Comprehensive Training Dynamics
We present the training loss and validation results for two distinct models, Llama3.1-8B and Qwen2.5-7B, in Fig 20. It can be observed that across different models, the training process of DaaR method consistently ensures gradual convergence, achieving high domain predictability and calibrated entropy fitting.

### v3

@@ D.2 Benchmarks
D.3 Training and Evaluation Details
@@
We implement our approaches using PyTorch [36] v2.4.1, coupled with PEFT v0.12.0 and the Transformers library [47] v4.45.2. Experiments are conducted on a computing platform equipped with four NVIDIA A100 GPUs (40GB), with pre-trained LLMs loaded as 16-bit floating-point numbers. The specific data-model development processes are completed in Data-Juicer Sandbox [9, 8], via integration with the ms-swift [61] training repository, and the OpenCompass [16] evaluation repository.
@@
Appendix E Details of Baselines
E.1 Implementation of Baselines
@@
Random Selection(Rand) A random selection of 8,000 data samples from the data pool was made, which to some extent reflects the distribution characteristics of the original data pool.
Instruction Length (IL) The length of the instruction can be considered a measure of input complexity. It is widely believed [5, 60] that more complex data is beneficial for enhancing model capabilities. Therefore, we select a subset of 8,000 entries with the maximum number of words (based on spaces) in the concatenation of Instruction and Input as part of the filtering process.
AlpaGasus [10] Using prompts to directly score and annotate the quality of the data leverages the powerful cognitive abilities of LLMs for evaluation and selection. Based on the original work, we use the GPT-3.5-Turbo API to score the data with the following prompt:
System Prompt:
@@
Instag [34] first utilizes ChatGPT to tag the samples based on semantics and intentions, then trains a LLaMA-based tagger on the ChatGPT tags to tag data. They use the number of tags as a proxy for complexity. We directly use ChatGPT-3.5-Turbo as a tagger to achieve better performance. Following the original paper, the prompt is as follows.
System Prompt:
@@
SuperFilter [29] introduces a method that astonishingly utilizes a small GPT2 model to successfully filter out the high-quality subset from the existing GPT4-generated instruction tuning dataset. The core concept is the instruction-following difficulty (IFD) score. By meticulously following this methodology, we utilize GPT-2 to select 8,000 data samples.
Deita [33] is an open-sourced project designed to facilitate Automatic Data Selection for instruction tuning in LLMs. It delves into the relationships between data complexity, quality, and diversity, and develops a series of methodologies. Specifically, we utilize it as the following three baselines:
@@
E.2 Visualization of Baseline’s Selected Data
@@ F.2 Validation on Qwen3-8B
Following the identical experimental setup outlined in Appendix D, we conducted experiments on Qwen3, with the results presented in Table 10. As demonstrated, DaaR achieved the top comprehensive score on Qwen3, surpassing all baselines by up to 4.7%. Notably, its performance on the MATH benchmark was particularly outstanding, exceeding the average of baselines by 25.91%. The comparative experiments affirm that DaaR maintains competitive effectiveness even on the SOTA LLM Qwen3-8B.
@@ F.3 Customized Data Selection Capability of DaaR
F.4 Robustness to Labeling Schemes and Embedding Representations
A core principle of DaaR is its "model-aware" design, which relies on the model’s own representations (via its embedding layer) and internal data structure (via clustering-based pseudo-labels) rather than external annotations or models. This section presents two ablation studies that investigate the validity and effectiveness of these design choices.
Analysis 1: Model-Aware Pseudo-Labels vs. Ground-Truth Labels
To analyze the impact of our clustering-based pseudo-labels, we conducted an experiment where we replaced them with human-annotated, ground-truth (GT) domain labels. This allowed us to directly compare a "model-aware" perspective against a "human-aware" one within our framework.
Setup. We reran the entire DaaR pipeline, using the GT labels to train the domain discrimination probe and subsequently to compute the predictive entropy for data selection. All other components remained identical.
Observations.
  • Data Selection Divergence: The choice of labels significantly altered the data selection process. As shown in Table 12, the overlap between the data selected using pseudo-labels and GT labels was only around 83-87%, indicating that they prioritize different data subsets.
    Table 12: Overlap of selected data between using pseudo-labels and ground-truth (GT) labels.
    Model Overlap (Pseudo vs. GT)
    Llama3.1-8B 87.3%
    Qwen2-7B 83.1%
    Qwen2.5-7B 84.9%
  • Final Performance: Critically, using GT labels led to a consistent, albeit small, decrease in average end-task performance across all models, as detailed in Table 13. This counter-intuitive result suggests that for fine-tuning, an LLM’s internal perception of data domains (captured by pseudo-labels) can be a more effective guide than human-defined categories. The GT labels may introduce biases that, while semantically correct to a human, are suboptimal for optimizing the model’s capabilities.
Table 13: Performance comparison: DaaR using model-aware pseudo-labels vs. GT labels. The "Diff" row shows the performance change when switching from pseudo-labels to GT labels.
Model Setting nq triviaqa hellaswag gsm8k math mbpp humaneval Avg
Llama3.1-8B DaaR w/ Pseudo 20.08 64.55 74.88 54.80 15.30 4.70 37.50 38.83
DaaR w/ GT 22.54 65.52 73.43 53.80 14.35 4.40 36.50 38.65
Diff +2.46 +0.97 -1.45 -1.00 -0.95 -0.30 -1.00 -0.18
Qwen2-7B DaaR w/ Pseudo 16.88 57.58 73.03 75.40 38.10 52.00 64.94 53.99
DaaR w/ GT 18.75 58.35 72.02 73.40 35.75 51.50 64.01 53.40
Diff +1.87 +0.77 -1.01 -2.00 -2.35 -0.50 -0.93 -0.59
Qwen2.5-7B DaaR w/ Pseudo 15.83 58.65 72.48 80.20 16.70 64.20 68.29 53.76
DaaR w/ GT 16.04 58.75 71.87 79.40 16.30 64.25 67.76 53.48
Diff +0.21 +0.10 -0.61 -0.80 -0.40 +0.05 -0.53 -0.28
Analysis 2: Internal vs. External Embedding
We also investigated the choice of using the model’s own embedding layer versus relying on powerful, external SOTA embedding models. While external models could provide an alternative "semantic space," we evaluated the trade-offs.
Setup. We processed our dataset using two SOTA external embedding models (GTE-Qwen2-7B and Qwen3-8B-Embedding) and compared the time cost against using the native embedding layer of the base LLM (Qwen2-7B).
Observations. While the external models were also capable of producing separable domain clusters, they introduced prohibitive computational overhead. As shown in Table 14, using external models was 175x to 192x slower and led to out-of-memory (OOM) errors on a 40GB GPU. This highlights the significant efficiency advantage of DaaR’s internal approach.
Table 14: Time cost for extracting embeddings from a 40K dataset.
Model/Layer Time Cost (Relative Speed)
Qwen2-7B (Internal Embedding Layer) 9.23s (1x)
GTE-Qwen2-7B (External Model) 1662s (175x slower)
Qwen3-8B-Embedding (External Model) 1774s (192x slower)
F.5 Analysis of End-to-End Stability
Given that DaaR is a multi-stage pipeline, it is important to analyze its end-to-end stability and robustness against potential compounding errors. To this end, we conducted three fully independent, end-to-end experimental runs of our entire framework and evaluated the consistency at critical stages of the pipeline. We present results for Qwen2-7B as a representative case.
Stability of Centroid Generation
The initial stage of our pipeline involves generating domain centroids. We measured the pairwise cosine similarities of the final centroid embeddings across the three independent runs. As shown in Table 15, the resulting centroids are remarkably consistent, with all similarity scores exceeding 0.98. This indicates that the foundational stage of our method is highly stable and is not a significant source of variance.
Table 15: Pairwise cosine similarity of domain centroids across three independent runs.
Domain Sim(Run 1-2) Sim(Run 1-3) Sim(Run 2-3)
Common Sense 0.9912 0.9895 0.9921
Reasoning 0.9845 0.9911 0.9887
Mathematics 0.9880 0.9803 0.9854
Coding 0.9905 0.9853 0.9918
Stability of Final Data Selection
We then examined whether minor variations in the initial stages propagate to the data selection outcome. We calculated the overlap ratio of the top 20% (8,000) selected samples across the three runs. Table 16 shows an average overlap of 96.4%, confirming that the final data selection is highly robust and largely deterministic, with minimal amplification of randomness.
Table 16: Overlap ratio of selected data subsets across three independent runs.
Metric Overlap (Run 1-2) Overlap (Run 1-3) Overlap (Run 2-3)
Overlap Ratio 95.7% 97.3% 96.1%
Consistency of Final Performance
Finally, the end-to-end stability of the framework is reflected in the final benchmark performance. As reported in Table 17, the results across the three runs are highly consistent, with a minimal standard deviation of 0.08 in the average score. This provides strong empirical evidence that our framework is robust and its outcomes are reliable, without significant compounding error effects.
Table 17: End-task performance across three independent runs on Qwen2-7B. Scores are reported with mean and standard deviation.
Run nq triviaqa hellaswag gsm8k math mbpp humaneval Avg
Run 1 17.92 56.78 72.91 75.00 39.60 51.40 64.80 54.06
Run 2 15.84 58.38 73.14 75.80 36.60 52.60 65.07 53.92
Run 3 15.18 57.96 72.99 76.70 38.80 51.40 65.24 54.04
Avg 16.31 57.71 73.01 75.83 38.33 51.80 65.04 54.01
Std. Dev. (1.43) (0.83) (0.12) (0.85) (1.55) (0.69) (0.22) (0.08)
F.6 Rationale for Entropy over Diversity as the Reward Signal
@@ G.6 Ablation of the Diversity Threshold
G.7 Ablation of Model Scales
To assess the generalizability of DaaR beyond the 7-8B parameter range, we conducted additional experiments on models of different scales: a smaller model, Llama3.2-3B, and a larger model, Qwen2.5-14B. Due to computational constraints, we performed a targeted comparison of DaaR against two key baselines: Raw (using the original, unselected data) and Random (a strong baseline involving random data selection).
The results are presented in Table 23. On the larger Qwen2.5-14B model, DaaR achieves the highest average score (61.54), outperforming the strong Random baseline. Similarly, on the smaller Llama3.2-3B model, DaaR (29.64) maintains a consistent advantage over both Raw and Random baselines.
These findings provide evidence that the effectiveness of DaaR is not confined to a specific model size and generalizes to both smaller and larger language models.
Table 23: Performance comparison on Llama3.2-3B and Qwen2.5-14B, demonstrating the generalization capability of DaaR across different model scales.
Model Method NQ TriviaQA Hellaswag GSM8K MATH MBPP HumanEval Avg
Qwen2.5-14B Raw 10.89 66.56 76.86 86.00 19.70 5.40 78.66 49.15
Random 20.06 65.80 76.76 86.40 36.90 64.20 75.46 60.80
DaaR 19.73 66.26 77.19 85.00 39.10 66.80 76.69 61.54
LLaMA3.2-3B Raw 7.62 53.10 68.77 26.60 4.50 3.80 23.17 26.79
Random 16.03 53.56 68.46 28.70 6.15 4.65 29.12 29.52
DaaR 15.13 53.79 68.23 29.10 5.63 5.30 30.34 29.64
G.8 Ablation of Downstream Sample Injection in Seed Generation
The seed generation process described in Section 4.1 includes a minor injection of downstream task samples. This component is intended to provide an initial diversity signal to accelerate the subsequent augmentation phase. To rigorously evaluate its impact, we present an ablation study comparing our standard method against a variant where this injection is completely removed.
It is important to note that in all settings, our experimental protocol maintains strict data partitioning. The downstream samples used for seeding, when present, are drawn from a set entirely disjoint from the datasets used for final evaluation.
Our analysis focuses on three key aspects: the impact on the final centroid representations, the effect on end-task model performance, and the change in data generation efficiency.
Impact on Centroid Similarity
We measured the cosine similarity between domain centroids generated with our standard method (w/ injection) and the ablation setting (w/o injection). The results in Table 24 show that the centroids are nearly identical, with an average similarity score of 0.987. This indicates that the minor injection has a negligible impact on the final learned domain representations.
Table 24: Cosine similarity of domain centroids generated with and without downstream seed injection, evaluated on Qwen2-7B.
Domain Similarity (w/ vs. w/o Injection)
Common Sense 0.9922
Reasoning 0.9865
Mathematics 0.9843
Coding 0.9878
Average 0.9877
Impact on Final Performance
As shown in Table 25, the end-task performance of models trained using data selected by DaaR is statistically indistinguishable between the two settings. This demonstrates that the effectiveness of our method is not dependent on the seed injection.
Table 25: End-task performance comparison between DaaR with and without downstream seed injection. The results show statistically insignificant differences.
Model Setting nq triviaqa hellaswag gsm8k math mbpp humaneval Avg
Llama3.1-8B DaaR w/ Inject 20.08 64.55 74.88 54.80 15.30 4.70 37.50 38.83
DaaR w/o Inject 20.39 64.80 76.05 55.40 13.65 5.75 36.48 38.93
Qwen2-7B DaaR w/ Inject 16.88 57.58 73.03 75.40 38.10 52.00 64.94 53.99
DaaR w/o Inject 16.22 58.33 73.41 75.20 36.37 52.95 65.14 53.95
Qwen2.5-7B DaaR w/ Inject 15.83 58.65 72.48 80.20 16.70 64.20 68.29 53.76
DaaR w/o Inject 14.91 58.32 72.55 79.70 16.30 63.70 70.65 53.73
Impact on Generation Efficiency
While not critical for performance, the injection significantly improves the efficiency of the diversity augmentation process. As detailed in Table 26, removing the initial seed diversity required approximately 4 times more generation attempts to meet the same diversity threshold ($\tau$). This confirms its functional role as a practical accelerator.
Table 26: Number of generation attempts required to meet the diversity threshold during seed augmentation.
Domain Attempts (w/ Injection) Attempts (w/o Injection)
Common Sense 31 134
Reasoning 61 248
Mathematics 236 819
Coding 21 105
G.9 Cost-Efficiency and Flexibility
@@ Appendix H Comprehensive Visualization and Experimental Results
Figure 17: The t-SNE visualization of embeddings for data samples on Qwen2-7B.
Figure 18: The t-SNE visualization of embeddings for data samples on Qwen2.5-7B.
@@ H.2 Visualization of Inter- & Intra-Diversity Distribution Data
Using the Qwen2-7B model as an example, we construct data based on the Inter-Diversity and Intra-Diversity distribution methods, selecting a batch of data every 20%. The visualization process is shown in Fig. 19 and Fig. 20. As seen in the figures, the data gradually transitions from domain-aware diverse to domain-aware closed, indicating that our data construction method effectively controls the distribution of different data.
@@ H.3 Comprehensive Training Dynamics
We present the training and validation dynamics for Llama3.1-8B and Qwen2.5-7B, in Fig 21. It can be observed that across different models, the training process of DaaR method consistently ensures gradual convergence, achieving high domain predictability and calibrated entropy fitting.